If it requires 5.45 s for a given sample of oxygen to effuse from a vessel at 300 K and 12.5 s for a similar size sample of an unknown gas to effuse from the same vessel at the same temperature, what is the experimentally determined molar mass of the unknown gas?
Answer #1

Grahams laws of differsion, Page No.: obortional to gas is inversel MA rate - Frolar mass - volume of gan equal time volume tt₂ = 5-45 s m = 32g/mol ti = 12.5t M = 2 [mol mass of al 5-45 2 12.5 12.5 mi (0.436 ) = 32 - 0.19 = 32 M My = 32 0.19 Molar m